# Efficient and accurate force replay in cosmological-baryonic simulations

### v1

@@ 3.1.1 Phase-space dependence at the final timestep
Figure 5: Preservation of phase-space distribution. Initial sample of all the selected stars (leftmost column), their distribution in the live simulation at “present day” (second from left), and the reconstructed distribution using the TEMP model for orbit integration (third from left). Distributions are shown in the XZ plane (top row) and for galactocentric distance vs.radial velocity (bottom row). The rightmost column show the median value of error metric for the initial distribution of sample stars in each bin. Most orbits have errors below 15%, with no angular dependence in the error metric. Errors are larger for particles closer to the center, deep in the disk potential.
@@
Furthermore, it is important to note the presence of a small satellite in our sample, approximately 40 kpc away with a radial velocity of 100 km $\textrm{s}^{-1}$. This satellite, while not associated with any subhalos in the halo catalog, appears to have bound stars in the sample. Additionally, an unbound stellar stream with a small bound progenitor is identified in our sample. As the simulation progresses, the satellite undergoes tidal disruption and phase mixing, resulting in a median orbit error in our reconstruction of roughly order 1, attributed to the TEMP model’s ignorance of self-gravity in the system. However, high-fidelity orbit reconstruction is observed for streams with errors below 5% along the stream track, except for the bound progenitor, which again exhibits an error of order 1. This emphasizes the critical role of self-gravity and the stripping times of stars. Ignoring these factors results in biased orbits.
@@
Fig. 6 plots $\frac{\Delta r}{r}$ at the final time step of integration as a function of integration time in periods passed for each orbit, color-coded by the true average distance from the host during the integration time. Also, the bound stars are shown with gray markers and majority of them show large errors. Approximately 70% of orbits retain phase-space information after the integration time, with $\frac{\Delta r}{r}\leq 1$ Comparing exact errors, only 13% of the recovered orbits have errors less than 10% (see Table. 3 for more statistics). Most of the orbits with higher relative errors ($\frac{\Delta r}{r}\geq 1$) are integrated for a prolonged period (over 10 periods) and are situated close to the center of the host (within 15 kpc).
@@ 4 Simulating stream formation
We use our TEMP model to reconstruct an example stellar stream with $\textrm{M}_{\star}=10^{7}$ Mdirect-$\odot$ at present day formed by tidal disruption of a dwarf satellite, identified by Panithanpaisal et al. (2021) in m12i. We begin tracking tidal disruption at the “stream formation time” for this progenitor galaxy, defined as in Panithanpaisal et al. (2021) to be the first time that the tidal deformation of the progenitor, measured by its principal axis ratio, exceeds a threshold value. For our example progenitor this occurs $6.5$ Gyr after the Big Bang, which we will call $T^{\prime}=0$ Gyr, representing the start of orbit integration. We assign unique stripping times ($T_{\textrm{strip}}$) to each star associated with the satellite by tracking when it crosses twice the time-evolving virial radius of the progenitor satellite: $|\vec{r}_{*}-\vec{r}_{\mathrm{sat}}|>2R_{200m}^{\mathrm{prog}}(t)$. Once a star reaches this distance, we consider it to be unaffected by the self gravity of its progenitor galaxy, which is not included in the TEMP model. We then begin orbit integration of the stripped star in the host potential up to the present day. The duration of orbit integration differs for each star, but does not exceed 7.2 Gyr for those stars that are assigned as stripped at or before $T^{\prime}=0$ Gyr. Once we have our potential models, these integrations are computationally inexpensive, allowing us to resimulate tidal stream formation at high resolution in a few minutes, without the cost of re-running the entire simulation.

### v2

@@ 4.1.1 Phase-space dependence at the final timestep
Figure 5: Preservation of phase-space distribution. Initial sample of all the selected stars (leftmost column), their distribution in the live simulation at “present day” (second from left), and the reconstructed distribution using the TEMP model for orbit integration (third from left). Distributions are shown in the XZ plane (top row) and for galactocentric distance vs.radial velocity (bottom row). The rightmost column shows the median value of error metric for the initial distribution of sample stars in each bin. Most orbits have errors below 15%, with no angular dependence in the error metric. Errors are larger for particles closer to the center, deep in the disk potential.
@@
We also note the presence of stars from a small satellite galaxy in our sample, approximately 40 kpc away with a radial velocity of 100 km $\textrm{s}^{-1}$. This satellite, while not associated with any subhalos in the halo catalog, appears to have bound stars in the sample. Our sample also contains stars from an unbound stellar stream with a small bound progenitor is identified in our sample (see Fig. 5). As the simulation progresses, the satellite undergoes tidal disruption and phase mixing, resulting in a median orbit error in our reconstruction of roughly order 1, attributed to the TEMP model’s ignorance of self-gravity in the system. However, high-fidelity orbit reconstruction is observed for streams with errors below 5% along the stream track, except for the bound progenitor, which again exhibits an error of order 1. This emphasizes the critical role of self-gravity and the stripping times of stars. Ignoring these factors results in biased orbits.
4.1.2 Positional errors at the final timestep
@@
Fig. 6 plots $\frac{\Delta r}{r}$ at the final time step of integration as a function of integration time in periods passed for each orbit, color-coded by the true average distance from the host during the integration time. Also, the bound stars are shown with gray markers and majority of them show large errors. Approximately 70% of orbits retain phase-space information after the integration time, with $\frac{\Delta r}{r}\leq 1$ Comparing exact errors, only 13% of the recovered orbits have errors less than 10% (see Table. 3 for more statistics).
Most of the orbits with higher relative errors ($\frac{\Delta r}{r}\geq 1$) are integrated for many dynamical times (over 10 periods) and are situated close to the center of the host (within 15 kpc). This is consistent with the fact that the largest fractional errors are in the central region (see Fig. 1). Additionally, these orbits frequently cross regions with stronger gravitational gradients, which can exacerbate the accumulation of integration errors.
Some stars on orbits with larger average galactocentric distances ($\sim 100$ kpc), also have larger errors. There are a few reasons for this. First, the orbit maybe highly eccentric, in which case the average distance may not accurately represent the orbit’s proximity to the center over time. Second, the phase sensitivity of the error metric amplifies errors near pericenter, leading to higher $\frac{\Delta r}{r}$ values, also matters most for highly radial orbits. Lastly, the omission of smaller-scale structures, such as subhalos, in our model may contribute to the discrepancies observed, as these can cause additional perturbations to the star particle in the simulation not accounted for in the TEMP model.
@@ 5 Simulating stream formation
We use our TEMP model to reconstruct an example stellar stream with $\textrm{M}_{\star}=10^{7}$ Mdirect-$\odot$ at present day formed by tidal disruption of a dwarf satellite of similar mass ($\textrm{M}_{\star}=10^{7}$ Mdirect-$\odot$), identified by Panithanpaisal et al. (2021) in m12i. The progenitor dwarf galaxy and subsequently the stream form self-consistently inside our cosmological zoom-in box m12i. This setup includes all relevant physical processes that influence stream formation and evolution, such as internal dynamical evolution, galactic tidal forces, and dynamical friction, all modeled at the resolution allowed by our particle fidelity. The resulting stream consists of around 2000 star particles. We begin tracking tidal disruption at the “stream formation time” for this progenitor galaxy, defined as in Panithanpaisal et al. (2021), which marks the first instance when the tidal deformation of the progenitor–measured by its principal axis ratio of the moment-of-inertia tensor–exceeds a threshold value. Since the moment-of-inertia tensor represents the overall shape of the progenitor, our disruption metric effectively captures shape deformation. Consequently, a small fraction of star particles ($\leq 0.5\%$) may already be unbound by the time we define stream formation. However, we include them for consistency in the analysis. For our example progenitor this occurs $6.5$ Gyr after the Big Bang, which we will call $T^{\prime}=0$ Gyr, representing the start of orbit integration. We assign unique stripping times ($T_{\textrm{strip}}$) to each star associated with the satellite by tracking when it crosses twice the time-evolving virial radius of the progenitor satellite: $|\vec{r}_{*}-\vec{r}_{\mathrm{sat}}|>2R_{200m}^{\mathrm{prog}}(t)$. Once a star reaches this distance, we consider it to be unaffected by the self-gravity of its progenitor galaxy, which is not included in the TEMP model. We then begin orbit integration of the stripped star in the host potential up to the present day. The duration of orbit integration differs for each star, but does not exceed 7.2 Gyr for those stars that are assigned as stripped at or before $T^{\prime}=0$ Gyr. Once we have our potential models, these integrations are computationally inexpensive, allowing us to resimulate tidal stream formation at high resolution in a few minutes, without the cost of re-running the entire simulation.